# Detecting Generative Parroting through Overfitting Masked Autoencoders

### v1

@@ 4 Experiments
WD AUG p_mask (%) Epochs Detection rate (%) $D_{\text{nov}}$ pass (%)
@@ 5 Conclusion
We presented a new approach for detecting generative parroting using an overfitted Masked Autoencoder with a Vision Transformer architecture. Our experiments demonstrate that training duration and model parameters, particularly the percentage mask (p_mask), play significant roles in the model’s ability to discern between seen, modified, and novel samples. The careful calibration of the loss threshold emerges as a crucial factor in mitigating the incidence of false positives, especially in large datasets where the distinction between original and parroted content is nuanced. Through extensive training, we observed that while a longer training time generally leads to higher detection rates for training and modified samples, it also increases the likelihood of incorrectly flagging novel samples as parroted. Our results underline the importance of selecting model configurations that balance sensitivity with specificity, thereby minimizing false positives without sacrificing detection accuracy.

### v3

@@ 4 Experiments
Method Num bins WD AUG p_mask (%) Epochs Detection rate (%) $D_{\text{nov}}$ pass (%)
@@
4.1 Benchmark
We compare our parroting detection algorithm against the graph hashing approaches which are commonly used for evaluating the novelty of generated CAD models [18, 6, 19]. The sketch is converted into a graph with edges representing curves and nodes representing curve end points and the centers of circles and arcs. The sketch is normalized and all point coordinates are quantized into equal sized bins. The quantized point coordinate values are then applied to the nodes as labels while edges are labelled using the curve type. The Weisfeiler-Lehman graph hash [15] can then be computed. This procedure assigns different hash values to sketches with different topology. Sketches with the same topology but small changes in geometry will receive the same graph providing the point coordinate values remain the same following quantization. Consequently the number of bins used for quantization provides control over the amount by which the geometry can change while preserving the same hash. In Table 1 we show results for 3, 7, 9, 31 and 63 bins.
4.2 Discussion
@@
The MAE parroting detector outperforms the detection rate of the graph based hashing benchmark for the $D_{\text{mod 1}}$ and $D_{\text{mod 2}}$ datasets, even when point coordinates are quantized into a 3 by 3 gird. Using this very small number of bins allows very large changes to the sketch point positions, casing the hash to be largely based on the sketch topology. As the edited sketches in our CAD variations dataset will have the same topology as the originals, graph based hashing approaches can unfairly exploit this isomorphism. Real generated results will tend to vary from the dataset with both small changes in sketch topology and geometry being present. Consequently the ability of the MAE parroting detector to achieve superior detection performance, with a comparable $D_{\text{nov}}$ pass percentage, and no strong dependence on sketch topology shows the utility of this technique for CAD parroting detection.
@@ 5 Conclusion
We presented a new approach for detecting generative parroting using an overfitted Masked Autoencoder with a Vision Transformer architecture. Our experiments demonstrate that training duration and model parameters, particularly the percentage mask (p_mask), play significant roles in the model’s ability to discern between seen, modified, and novel samples. The careful calibration of the loss threshold emerges as a crucial factor in mitigating the incidence of false positives, especially in large datasets where the distinction between original and parroted content is nuanced. Through extensive training, we observed that while a longer training time generally leads to higher detection rates for training and modified samples, it also increases the likelihood of incorrectly flagging novel samples as parroted. Our results underline the importance of selecting model configurations that balance sensitivity with specificity, thereby minimizing false positives without sacrificing detection accuracy. The approach is shown to achieve superior performance to the hash based parroting detection algorithms employed in previous works, while not explicitly exploiting the isomorphic sketch graphs created when sketches are varied parametrically.
@@
Acknowledgments
We would like to thank Ali Mahdavi-Amiri from Simon Fraser University and Karl D.D. Willis from Autodesk Research for the initial discussions on this work, which provided valuable insights.